What is the maximum height of h=96t-16t power of two?
Answer by Alan3354(69443)

What is the maximum height of h=96t-16t power of two?
-------------------
It's at the vertex of the parabola.
That's on the line of symmetry, t = -b/2a = -96/-32
h(3) = 288 - 144
= 144 feet max height